A box contains six ham and four cheese sandwiches. Two sandwiches are randomly selected and eaten. What is the probability that both are ham sandwiches?
Found 2 solutions by stanbon, scott8148:
Answer by stanbon(75887)

A box contains six ham and four cheese sandwiches. Two sandwiches are randomly selected and eaten. What is the probability that both are ham sandwiches?
---
# of pairs of ham sandwiches: 6C2 = 15
---
# of pairs of samdwiches: 10C2 = 45
----
P(select 2 ham) = 15/45 = 1/3
=================================
Cheers,
Stan H.
===============

Answer by scott8148(6628)

there are 10C2 ways to select 2 sandwiches

there are 6C2 ways to select 2 HAM sandwiches

so the probability of selecting two ham sandwiches is ___(6C2) / (10C2)